Statistical Analysis of Pile Games using Symbolic Computation Thotsaporn “Aek” THANATIPANONDA and Doron ZEILBERGER

Abstract. The Maple packages. This article is accompanied by three Maple packages

• SnakesAndLadders.txt
• PosPileGames.txt
• GenPileGames.txt

They are available, along with numerous input and output files, from the front of this article

http://www.math.rutgers.edu/~zeilberg/mamarim/mamarimhtml/pile.html .

References

[LT] Ho-Hon Leung and Thotsaporn ”Aek” Thanatipanonda, A Probabilistic Two-Pile Game, to

appear. https://arxiv.org/abs/1903.03274.

[MM] David Mccune and Lori Mccune, Counting your chickens with Markov chains, Mathematics Magazine 92 (2019), 162-172.

[T] Thotsaporn ”Aek” Thanatipanonda, A Quantitative Study on Average Number of Spins of

Two-Player Dreidel, https://arxiv.org/abs/1907.11851 .

Thotsaporn “Aek” Thanatipanonda, Mahidol University International College, Nakornpathom, Thailand

Email: thotsaporn at gmail dot com .

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: DoronZeil at gmail dot com .

Written: July 18, 2019.